find the equation of the line with slope 8 and passing through(-5,-6).....
Answer by Alan3354(69443)

find the equation of the line with slope 8 and passing through(-5,-6).....
-------------------------
OK, I will.
y-y1 = m*(x-x1) (x1,y1) is (-5,-6), m is the slope
y+6 = 8*(x+5)
y+6 = 8x+40
y = 8x+34 (slope-intercept form)
8x - y = -34 (standard form)